# Model of incompressible turbulent flows via a kinetic theory

### v1

@@ 4 Extension to wall-bounded turbulence
The kinetic turbulence model developed above is valid only for fully turbulent flows, which we refer to as the “high-Reynolds-number kinetic (HR-BGK) model”. However, in the near-wall region for wall-bounded flows, the turbulence is strongly damped and the effects of molecular viscosity become paramount (hanjalic1976contribution). Furthermore, the turbulent motion is restricted and the relaxation time $\tau$ and the turbulent Reynolds number $\text{Re}_{T}=K^{2}/(\epsilon\nu_{0})$ tend to zero. To address this challenge, a natural approach is to employ wall function methods from conventional turbulence modelling, which apply the HR-BGK model outside the viscous sublayer using empirical near-wall correlations. Alternatively, to enable resolution of flows within the viscous sublayer, we develop a low-Reynolds-number kinetic (LR-BGK) model that incorporates damping functions and an additional viscous diffusion term.
In the following subsections, we present the LR-BGK model, and the wall boundary conditions including non-slip condition for the LR-BGK model and wall function condition for the HR-BGK model.
@@ 4.1 Low-Reynolds-number kinetic model
$\tau=\frac{f_{\tau}}{7}\frac{K}{\epsilon},$ (36)
where $f_{\tau}$ is a damping function corresponding to $f_{\mu}$. To match the turbulent viscosity definition in (23) with the classical expression incorporating the van Driest damping function $f_{\mu}$, we equate:
@@
Then with the relationship $K^{eq}=K-\tau\epsilon$ and the expression (36), we can obtain
$\frac{2f_{\tau}\left(1-f_{\tau}/7\right)}{21}=C_{\mu}f_{\mu},$ (38)
which gives
$f_{\tau}=\frac{7-\sqrt{49-24f_{\mu}}}{2}.$ (39)
Note that here the physical constraint $0\leq f_{\tau}\leq 1$ is considered to ensure physically meaningful values. A number of damping functions have been proposed. For instance, the expression proposed by jones1972prediction depends solely on $\text{Re}_{T}$. Subsequent refinements by myong1990new and nagano1990improved introduce dependencies on the dimensionless wall coordinate ($y^{+}$), thus explicitly incorporating sensitivity to near-wall length scales into the turbulence model.
@@
$\partial_{t}F+\bm{\xi}\cdot\nabla F+\bar{\bm{a}}\cdot\nabla_{\bm{\xi}}F=\mathcal{C}(F)+\mathcal{S},$ (40)
@@
Alternatively, following the well-known Shakhov model (shakhov1968generalization) for non-equilibrium rarefied gas flows and applying the Hermite‐polynomial expansion method, we can obtain an alternative formulation,
$\mathcal{S}=\frac{3\nu_{0}\nabla^{2}K}{2K^{eq}}\left(\frac{\left(\bm{\xi}-\bm{U}\right)^{2}}{2K^{eq}}-1\right)F^{eq}.$ (43)
Note that the above construction is formally equivalent to replacing $F$ with its equilibrium counterpart $F^{eq}$ in (42). This approximation is justified, because the viscous diffusion of TKE is relevant only when the turbulent Reynolds number $\text{Re}_{T}$ (and therefore the relaxation time $\tau$) is small, a regime in which the VDF remains close to equilibrium.
@@ 4.2 Wall boundary conditions
where $U_{\sigma}=\sqrt{\sigma_{w}}$ is the friction velocity obtained from the wall shear stress $\sigma_{w}$ with unity density, $\kappa\approx 0.41$ is the von Kármán constant, and $B$ is an empirical constant related to the thickness of the viscous sublayer; $y$ denotes the wall-normal distance measured from the wall into the fluid and $y^{+}=U_{\sigma}y/\nu_{0}$ is its dimensionless form. The corresponding wall value of the TKE is specified by the usual log-layer estimate,
$K_{w}=U^{2}_{\tau}/\sqrt{C_{\mu}}.$ (49)
@@
where $\bm{x}_{nc}$ is the neighboring cell of $\bm{x}_{w}$. The equilibrium term enforces the prescribed wall function moments, while the second term extrapolates the non-equilibrium part of the VDF, ensuring compatibility with the interior solution. It is worth noting that the non-equilibrium extrapolation method assumes that the deviation from equilibrium is small to permit linear extrapolation. For highly non-equilibrium turbulent flows, this method may fail to work properly.
@@ 5 Study of turbulent plane Couette flow
The turbulent plane Couette flow provides an ideal benchmark (henry1984analytical; nisizima1987turbulent; schneider1989reynolds; andersson1994modeling) to test whether the present kinetic model can accurately capture wall-bounded shear turbulence. The LR-BGK and HR-BGK models, coupled with the diffuse and non-equilibrium extrapolation boundary conditions, respectively, are employed to simulate this flow at different Reynolds numbers. The central issue in wall-bounded shear turbulence is the prediction of mean velocity profiles and the derivation of friction laws. In this study, we consider the configuration of steady turbulent Couette flow between two infinite parallel plates located at $y=\pm h$, as illustrated in figure 1. The two plates move in opposite directions with constant velocities $\pm U_{w}$ in the $x$-direction, respectively. The corresponding Reynolds number is defined as $\text{Re}=U_{w}h/\nu_{0}$.
Figure 1: Sketch of turbulent plane Couette flow between two infinite parallel plates separated by $2h$. The plates move in opposite directions with constant velocities $\pm U_{w}$ in the $x$-direction, generating a monotonic mean velocity profile $U_{x}(y)$.
@@ 5.1 Numerical method and parameter set-up
As mentioned in §2.2, the dissipation rate $\epsilon$ is an input parameter to determine the relaxation time $\tau$ in the present kinetic model. For the HR-BGK model, we adopt the widely used transport equation for the dissipation rate $\epsilon$ proposed by launder1983numerical to calculate $\epsilon$,
@@
where $\nu_{T}=0.09K^{2}/\epsilon$, $C_{\epsilon}=1.3$, $C_{\epsilon 1}=1.44$ and $C_{\epsilon 2}=1.92$. It is worth highlighting that both the turbulent kinetic energy $K$ and the stress tensor $\bm{\sigma}$ are obtained directly from the solution of kinetic equation; therefore, equation (51) retains higher-order non-equilibrium information that is absent from the linear eddy viscosity model. The corresponding wall boundary condition for the Couette flow employs a wall function approach, where the wall value of $\epsilon$ is specified according to the usual logarithmic layer estimate,
$\epsilon_{w}=\frac{U_{\sigma}^{3}}{\kappa y_{1}},$ (52)
with $y_{1}$ being the wall-normal distance from the wall to the center of the first grid cell adjacent to the wall.
On the other hand, for the LR-BGK model, an alternative transport equation for $\epsilon$ with wall damping effect (nagano1990improved) is adopted,
$\frac{\partial\epsilon}{\partial t}+\left(\bm{U}\cdot\nabla\right)\epsilon=\nabla\cdot\left(\frac{\nu_{T}}{C_{\epsilon}}\nabla\epsilon\right)+C_{\epsilon 1}\frac{\epsilon}{K}\bm{\sigma}:\nabla\bm{U}-C_{\epsilon 2}f_{\epsilon 2}\frac{\epsilon^{2}}{K},$ (53)
where $\nu_{T}=0.09f_{\mu}K^{2}/\epsilon$, $C_{\epsilon}=1.3$, $C_{\epsilon 1}=1.45$ and $C_{\epsilon 2}=1.92$. $f_{\mu}$ and $f_{\epsilon 2}$ are the damping functions, defined as
$f_{\mu}$ $=\left(1-\exp\left(-\frac{-y^{+}}{26}\right)\right)^{2}\left(1+\frac{4.1}{\text{Re}_{T}^{3/4}}\right),$ (54)
$f_{\epsilon 2}$ $=\left(1-0.3\exp{\left(-\frac{\text{Re}_{T}^{2}}{42.5}\right)}\right)\left(1-\exp\left(-\frac{-y^{+}}{6}\right)\right)^{2}.$
The corresponding wall boundary condition is
@@
with $a$ being a dimensionless stretching parameter controlling the characteristics of the mesh distribution, and $N_{y}$ is the number of mesh cells. In the present study, $a$ is taken as $2.5$. We utilize $N_{y}=401$ cells for the LR-BGK model, ensuring that the center of the first off-wall cell lies within the viscous sublayer for all cases considered. The velocity space truncated at $5\sqrt{U_{w}},5\sqrt{U_{w}}]$ is discretized by the trapezoidal rule (wu2014solving) with 4000 non-uniform velocity grid points. Furthermore, the time step is set to $0.05\times 5\sqrt{U_{w}}/(\Delta y_{min})$ with $y_{min}$ being the minimum cell size. The steady-state solution is considered to be achieved when the maximum relative change in the velocity field between two consecutive iterations falls below $10^{-10}$.
@@ 5.2 Comparison with experiments and direct numerical simulations
The mean velocity profiles predicted by the LR-BGK and HR-BGK models are compared with the experimental (bech1995investigation; robertson1959study) and DNS (bech1995investigation; pirozzoli2014turbulence) data at $\text{Re}=1300$ and $10133$, as shown in figures 2 and 3. The results are presented in both outer units (scaled by $U_{w}$ and $h$) and wall units (scaled by $U_{\sigma}$ and $y^{+}$) to examine the performance of the models across different flow regions. In the case of $\text{Re}=1300$, Figure 2a demonstrates that both models accurately capture the overall velocity distribution in outer coordinates, closely matching the experimental and DNS data of bech1995investigation. Specifically, the LR-BGK model shows excellent agreement throughout the entire flow domain, while the HR-BGK model exhibits slightly larger deviations in the near-wall region, which can be expected since the viscous sublayer is bypassed in the wall function boundary condition. At a higher Reynolds number ($\text{Re}=10133$), figure 3 shows that both the HR-BGK and LR-BGK models maintain their predictive accuracy, in good agreement with the experimental data of robertson1959study and DNS data of pirozzoli2014turbulence, though both models exhibit slightly larger deviations in the near-wall region. The logarithmic law profile (48) is well-captured by both models with $\kappa=0.41$, using $B=6.5$ and $7.0$ for $\text{Re}=1300$ and $10133$, respectively.
Figure 2: Mean velocity profiles scaled in (a) outer and (b) wall units of the turbulent Couette flow. The solid and dash‑dot lines represent the predictions of the LR-BGK model and the HR-BGK model at $\text{Re}=1300$, respectively. The diamonds denote the experimental data of bech1995investigation at $\text{Re}=1260$, and the circles denote the DNS data of bech1995investigation at $\text{Re}=1300$. The dashed line represents the classical logarithmic law profile $\left(\ln y^{+}\right)/0.41+6.5$.
Figure 3: Mean velocity profiles scaled in (a) outer and (b) wall units of turbulent Couette flow. The solid and dash‑dot lines represent the predictions of the LR-BGK model and the HR-BGK model at $\text{Re}=10133$, respectively. The diamonds denote the experimental data of robertson1959study at $\text{Re}=10000$, and the circles denote the DNS data of pirozzoli2014turbulence at $\text{Re}=10133$. The dashed line represents the classical logarithmic law profile $\left(\ln y^{+}\right)/0.41+7$.
@@
Figure 4: Variation of skin friction coefficient with Reynolds number. The solid and dash‑dot lines represent the predictions of the LR-BGK model and the HR-BGK model, respectively. Open symbols refer to experimental data by kitoh2005experimental (triangles), el1980velocity (diamonds), reichardt1956geschwindigkeitsverteilung (right‑pointing triangles), robertson1959turbulent (pluses). Filled symbols refer to DNS data by bech1995investigation (left‑pointing triangles), tsukahara2006dns (squares), pirozzoli2014turbulence (circles).
@@
The Reynolds shear stress profiles provide a stringent test of the capacity of the models to capture turbulent momentum transport mechanisms. Figure 5 compares the dimensionless Reynolds shear stress $\langle u^{\prime}_{x}u^{\prime}_{y}\rangle/\sigma_{w}$ predicted by both kinetic models with DNS data (bech1995investigation; pirozzoli2014turbulence) as Re ranges from 1300 to 10133. It is noteworthy that the HR-BGK model, due to its coarser grid resolution, exhibits minor numerical oscillations in the near-wall region. Both kinetic models exhibit excellent agreement with the DNS data, accurately capturing both the peak value and its location, particularly in the logarithmic and outer regions ($y^{+}\geq 15$). However, the LR-BGK model shows some deviations in the transition region ($1<y^{+}<15$) due to the inherent complexity of this region where viscous and turbulent effects compete.
@@
Figure 5: Reynolds shear stress profiles scaled in wall units of turbulent Couette flow at (a) $\text{Re}=1300$ and (b) $\text{Re}=10133$. The solid and dash‑dot lines represent the predictions of the LR-BGK model and the HR-BGK model, respectively. The circles and the diamonds denote the DNS data of bech1995investigation and pirozzoli2014turbulence, respectively.
@@
Figures 6 and 7 present detailed comparisons of the velocity fluctuation intensities (normal Reynolds stress) in all three spatial directions which can provide insight into the anisotropic nature of near-wall turbulence, which the linear eddy viscosity model (22) cannot capture (launder1974application). The velocity fluctuation intensities predicted by the present models show noticeable differences from the DNS data (bech1995investigation; pirozzoli2014turbulence), with the discrepancies becoming more pronounced at higher Re. At $\text{Re}=1300$, the dimensionless streamwise $\sqrt{\left\langle(u^{\prime}_{x})^{2}\right\rangle}$ and vertical $\sqrt{\left\langle(u^{\prime}_{z})^{2}\right\rangle}$ components exhibit relatively small deviations from DNS data near the channel centerline, while the differences become increasingly evident as Re increases to $10133$. The LR-BGK model demonstrates superior performance in capturing the peak locations of the streamwise fluctuations, where the peak occurs around $y^{+}\approx 10$. Although the present models do not provide a clear separation between the normal stress components, they demonstrate the capability to capture non-Newtonian effects (nonlinear eddy viscosity effects) as mentioned in §3. This non-equilibrium phenomenon, often reported in rarefied gas flows at finite Knudsen numbers, exhibits the anisotropic stress behaviour that cannot be captured by the NS equations (xu2007multiple).
@@
Figure 6: Distribution of the velocity fluctuation intensities (Reynolds normal stress) scaled in wall units of turbulent Couette flow from (a) the LR-BGK model and (b) the HR-BGK model at $\text{Re}=1300$. The dash, dash-dot and dot lines represent $\sqrt{\left\langle(u_{x}^{\prime})^{2}\right\rangle}/U_{\sigma}$, $\sqrt{\left\langle(u_{y}^{\prime})^{2}\right\rangle}/U_{\sigma}$ and $\sqrt{\left\langle(u_{z}^{\prime})^{2}\right\rangle}/U_{\sigma}$, respectively. The symbols denote the DNS data of bech1995investigation: $\sqrt{\left\langle(u_{x}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (circles), $\sqrt{\left\langle(u_{y}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (squares) and $\sqrt{\left\langle(u_{z}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (diamonds).
@@
Figure 7: Distribution of the velocity fluctuation intensities (Reynolds normal stress) scaled in wall units of turbulent Couette flow from (a) the LR-BGK model and (b) the HR-BGK model at $\text{Re}=10133$. The dash, dash-dot and dot lines represent $\sqrt{\left\langle(u_{x}^{\prime})^{2}\right\rangle}/U_{\sigma}$, $\sqrt{\left\langle(u_{y}^{\prime})^{2}\right\rangle}/U_{\sigma}$ and $\sqrt{\left\langle(u_{z}^{\prime})^{2}\right\rangle}/U_{\sigma}$, respectively. The symbols denote the DNS data of pirozzoli2014turbulence: $\sqrt{\left\langle(u_{x}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (circles), $\sqrt{\left\langle(u_{y}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (squares) and $\sqrt{\left\langle(u_{z}^{\prime})^{2}\right\rangle}/U_{\sigma}$ (diamonds).

### v2

@@ 4 Extension to wall-bounded turbulence
The kinetic turbulence model developed above is valid only for fully turbulent flows, which we refer to as the “high-Reynolds-number kinetic (HR-BGK) model”. However, in the near-wall region for wall-bounded flows, the turbulence is strongly damped and the effects of molecular viscosity become paramount (hanjalic1976contribution). Furthermore, the turbulent motion is restricted and the relaxation time $\tau$ and the turbulent Reynolds number $\text{Re}_{T}=K^{2}/(\epsilon\nu_{0})$ tend to zero. To address this challenge, a natural approach is to employ wall function methods from conventional turbulence modelling, which apply the HR-BGK model outside the viscous sublayer using empirical near-wall correlations. Alternatively, to enable resolution of flows within the viscous sublayer, we develop a low-Reynolds-number kinetic (LR-BGK) model that incorporates damping functions and an additional viscous diffusion term. In the following subsections, we present the LR-BGK model, and the wall boundary conditions including non-slip condition for the LR-BGK model and wall function condition for the HR-BGK model.
@@ 4.1 Low-Reynolds-number kinetic model
$\tau=\frac{f_{\tau}}{7}\frac{K}{\epsilon},\qquad f_{\tau}=\frac{7-\sqrt{49-24f_{\mu}}}{2},$ (36)
where $f_{\tau}$ is a damping function corresponding to $f_{\mu}$. Note that here the physical constraint $0\leq f_{\tau}\leq 1$ is considered to ensure physically meaningful values. Substituting this into the definition of turbulent viscosity (23) yields
@@
A number of damping functions have been proposed. For instance, the expression proposed by jones1972prediction depends solely on $\text{Re}_{T}$. Subsequent refinements by myong1990new and nagano1990improved introduce dependencies on the dimensionless wall coordinate ($y^{+}$), thus explicitly incorporating sensitivity to near-wall length scales into the turbulence model.
@@
$\partial_{t}F+\bm{\xi}\cdot\nabla F+\bar{\bm{a}}\cdot\nabla_{\bm{\xi}}F=\mathcal{C}_{\text{BGK}}(F)+\mathcal{S},$ (38)
@@ 4.2 Wall boundary conditions
where $U_{\sigma}=\sqrt{\sigma_{w}}$ is the friction velocity obtained from the wall shear stress $\sigma_{w}$ with unity density, $\kappa\approx 0.41$ is the von Kármán constant, and $B$ is an empirical constant; $y$ denotes the wall-normal distance measured from the wall into the fluid and $y^{+}=U_{\sigma}y/\nu_{0}$ is its dimensionless form. In addition, the corresponding wall value of the TKE is specified by the usual log-layer estimate, $K_{w}=U^{2}_{\sigma}/\sqrt{C_{\mu}}$.
@@
where $\bm{x}_{nc}$ is the neighboring cell of $\bm{x}_{w}$. The equilibrium term enforces the prescribed wall function moments, while the second term extrapolates the non-equilibrium part of the VDF, ensuring compatibility with the interior solution.
@@ 5 Study of turbulent plane Couette flow
The turbulent plane Couette flow provides an ideal benchmark (henry1984analytical; nisizima1987turbulent; schneider1989reynolds; andersson1994modeling) to test whether the present kinetic model can accurately capture wall-bounded shear turbulence. The LR-BGK and HR-BGK models, coupled with the diffuse and non-equilibrium extrapolation boundary conditions, respectively, are employed to simulate this flow at different Reynolds numbers. The central issue in wall-bounded shear turbulence is the prediction of mean velocity profiles and the derivation of friction laws. In this study, we consider the configuration of steady turbulent Couette flow between two infinite parallel plates located at $y=\pm h$. The two plates move in opposite directions with constant velocities $\pm U_{w}$ in the $x$-direction, respectively. The corresponding Reynolds number is defined as $\text{Re}=U_{w}h/\nu_{0}$.
@@ 5.1 Numerical method and parameter set-up
As mentioned in §2.2, the dissipation rate $\epsilon$ is an input parameter to determine the relaxation time $\tau$ in the present kinetic model. For the HR-BGK model, we adopt the widely used transport equation for the dissipation rate proposed by launder1983numerical to calculate $\epsilon$,
@@
where $\nu_{T}=0.09K^{2}/\epsilon$, $C_{\epsilon}=1.3$, $C_{\epsilon 1}=1.44$ and $C_{\epsilon 2}=1.92$. It is worth highlighting that both the turbulent kinetic energy $K$ and the stress tensor $\bm{\sigma}$ are obtained directly from the solution of kinetic equation; therefore, equation (47) retains higher-order non-equilibrium information that is absent from the linear eddy viscosity model. The wall boundary condition for Couette flow is implemented using a wall function approach. In this approach, the wall dissipation rate $\epsilon_{w}$ is prescribed based on the standard logarithmic layer estimate as $\epsilon_{w}=U_{\sigma}^{3}/(\kappa y_{1})$, where $y_{1}$ is the wall-normal distance from the wall to the center of the first grid cell adjacent to the wall.
Conversely, for the LR-BGK model, we employ an alternative $\epsilon$-transport equation developed by nagano1990improved, where the turbulent viscosity and model coefficients are modified as $\nu_{T}=0.09f_{\mu}K^{2}/\epsilon$, $C_{\epsilon 1}=1.45$ and $C_{\epsilon 2}=1.92f_{\epsilon 2}$. The damping functions $f_{\mu}$ and $f_{\epsilon 2}$ are utilized to account for near-wall damping, with their specific forms defined therein. The corresponding wall boundary condition is,
@@
with $a$ being a dimensionless stretching parameter controlling the characteristics of the mesh distribution, and $N_{y}$ is the number of mesh cells. In the present study, $a$ is taken as $2.5$. We utilize $N_{y}=401$ cells for the LR-BGK model, ensuring that the center of the first off-wall cell lies within the viscous sublayer for all cases considered. The velocity space truncated at $5\sqrt{U_{w}},5\sqrt{U_{w}}]$ is discretized by the trapezoidal rule (wu2014solving) with 4000 non-uniform velocity grid points. Furthermore, the time step is set to $0.05\times 5\sqrt{U_{w}}/(\Delta y_{min})$ with $\Delta y_{min}$ being the minimum cell size, which satisfies the Courant-Friedrichs-Lewy (CFL) condition. The steady-state solution is considered to be achieved when the maximum relative change in the velocity field between two consecutive iterations falls below $10^{-10}$.
@@ 5.2 Comparison with experiments and direct numerical simulations
The mean velocity profiles predicted by the LR-BGK and HR-BGK models are compared with the experimental (bech1995investigation; robertson1959study) and DNS (bech1995investigation; pirozzoli2014turbulence) data at $\text{Re}=1300$ and $10133$, as shown in figures 1 and 2. The results are presented in both outer units (scaled by $U_{w}$ and $h$) and wall units (scaled by $U_{\sigma}$ and $y^{+}$) to examine the performance of the models across different flow regions. In the case of $\text{Re}=1300$, Figure 1a demonstrates that both models accurately capture the overall velocity distribution in outer coordinates, closely matching the experimental and DNS data of bech1995investigation. Specifically, in figure 1b, the HR-BGK model shows excellent agreement throughout the entire flow domain, while the LR-BGK model exhibits slightly larger deviations in the logarithmic layer. This is because HR-BGK targets the outer region (bypassing the viscous sublayer), whereas LR-BGK includes stronger near-wall corrections that can slightly influence the log-layer profile. At a higher Reynolds number ($\text{Re}=10133$), figure 2 shows that both the HR-BGK and LR-BGK models maintain their predictive accuracy, in good agreement with the experimental data of robertson1959study and DNS data of pirozzoli2014turbulence, though both models exhibit slightly larger deviations in the near-wall region. The logarithmic law profile (45) is well-captured by both models with $\kappa=0.41$, using $B=6.5$ and $7.0$ for $\text{Re}=1300$ and $10133$, respectively.
Figure 1: Mean velocity profiles scaled in (a) outer and (b) wall units of the turbulent Couette flow at $\text{Re}=1300$.
Figure 2: Mean velocity profiles scaled in (a) outer and (b) wall units of turbulent Couette flow at $\text{Re}=10133$.
@@
Figure 3: Variation of skin friction coefficient with Reynolds number.
@@
Figure 4 compares the dimensionless Reynolds shear stress $\langle u^{\prime}_{x}u^{\prime}_{y}\rangle/\sigma_{w}$ predicted by both kinetic models with DNS data (bech1995investigation; pirozzoli2014turbulence) as Re ranges from 1300 to 10133. It is noteworthy that the HR-BGK model, due to its coarser grid resolution, exhibits minor numerical oscillations in the near-wall region. Both kinetic models exhibit excellent agreement with the DNS data, accurately capturing both the peak value and its location, particularly in the logarithmic and outer regions ($y^{+}\gtrsim 15$). However, the LR-BGK model shows some deviations in the transition region ($1\lesssim y^{+}\lesssim 15$) due to the inherent complexity of this region where viscous and turbulent effects compete.
@@
Figure 4: Reynolds shear stress profiles scaled in wall units of turbulent Couette flow at (a) $\text{Re}=1300$ and (b) $\text{Re}=10133$.
@@
Figures 5 and 6 present detailed comparisons of the velocity fluctuation intensities (normal Reynolds stress) in all three spatial directions which can provide insight into the anisotropic nature of near-wall turbulence, which the linear eddy viscosity model (22) cannot capture (launder1974application). The velocity fluctuation intensities predicted by the present models show noticeable differences from the DNS data (bech1995investigation; pirozzoli2014turbulence), with the discrepancies becoming more pronounced at higher Re. At $\text{Re}=1300$, the dimensionless streamwise $\sqrt{\left\langle(u^{\prime}_{x})^{2}\right\rangle}$ and vertical $\sqrt{\left\langle(u^{\prime}_{z})^{2}\right\rangle}$ components exhibit relatively small deviations from DNS data near the channel centerline, while the differences become increasingly evident as Re increases to $10133$. The LR-BGK model demonstrates superior performance in capturing the peak locations of the streamwise fluctuations, where the peak occurs around $y^{+}\approx 10$. Notably, the largest discrepancies relative to DNS occur in the near-wall/buffer region (approximately $y^{+}\approx 1\sim 10$), where the separation among the normal stress components is underpredicted; this highlights a limitation of the present formulation and motivates further improvements. Although the magnitudes of the normal stress components are consistent with DNS, a clear separation among these components is still not reproduced by the present models. Nevertheless, they demonstrate the capability to capture non-Newtonian (nonlinear eddy viscosity) effects as mentioned in §3. This non-equilibrium phenomenon, often reported in rarefied gas flows at finite Knudsen numbers, exhibits the anisotropic stress behaviour that cannot be captured by the NS equations (xu2007multiple).
@@
Figure 5: Distribution of the velocity fluctuation intensities (Reynolds normal stress) scaled in wall units of turbulent Couette flow from (a) the LR-BGK model and (b) the HR-BGK model at $\text{Re}=1300$.
@@
Figure 6: Distribution of the velocity fluctuation intensities (Reynolds normal stress) scaled in wall units of turbulent Couette flow from (a) the LR-BGK model and (b) the HR-BGK model at $\text{Re}=10133$.